OpenStudy (anonymous):

Operations with complex numbers? 7i*3i(-8-6i). HELP PLEASE!

ganeshie8 (ganeshie8):

i^2 = -1

ganeshie8 (ganeshie8):

so, 7i*3i = 21*i^2 = 21(-1) = -21

ganeshie8 (ganeshie8):

-21(-8-6i) distribute..

OpenStudy (anonymous):